OpenStudy (anonymous):

Find midoint (-3,3)(0,-3) what if i have a zero in one of the numerators?

OpenStudy (lgbasallote):

midpoint is \[x_m = \frac{x_1 + x_2}{2} \implies \frac{-3 + 0}{2}\]

OpenStudy (lgbasallote):

that's the x coordinate of the midpoint

OpenStudy (lgbasallote):

y coordinate of the midpoint is \[y_m = \frac{y_1 + y_2}{2} \implies \frac{3 + -3}{2}\]

OpenStudy (lgbasallote):

get it?

OpenStudy (anonymous):

yes but the y coordinate ends up being 0/2

OpenStudy (anonymous):

i cant have a zero in the numerator can i?

OpenStudy (lgbasallote):

0 divided by anything is 0 :)

OpenStudy (anonymous):

haha thanks, i think you have helped me before but thanks again!

OpenStudy (lgbasallote):

really? hmm i must be a hotshot here :O lol jk